a voltmeter v of 23 k ohm resistance in series with a resistor r across a 230v supply reads 92v. calculate the value of the resistor r 


1
Expert's answer
2021-05-06T17:06:03-0400


Given :

Resistance of Voltmeter"=R"V"=23k""\Omega"

Reading of voltmeter = Potential difference across the voltmeter "=92V"

External source is applied of voltage "=230V"

So Potential difference across resistance r will be "= 230V-92V= 138V"

To calculate the value of resistance r

from Diagram

Current through the voltmeter and resistance r will be


"I= \frac{230V}{R_v + r}" by Ohm's Law

Now potential difference across voltmeter = reading of voltmeter "=92V" "= IR_v" "= \frac{230R_V}{R_v+r}"

after putting value of "R"v

"\implies" "\frac{r}{R_v} = \frac{230}{92}-1"

"\implies r = 1.5 \times R_v" = "34500\Omega"

Final answer

The value of resistance r connected in series with voltmeter is "34.5k\Omega" .